OpenStudy (anonymous):

d/dx[ln((x-1)^3(x^2+1)^4)] I tried but could not get the right answer. I used the chain rule. The book says the answer is 3/(x-1) + 8x/x^2+1 how do I get rid of the powers I get the same but with powers at the bottom.

myininaya (myininaya):

so we want to differentiate \[y=\ln[(x-1)^3](x^2+1)^4 \]?

myininaya (myininaya):

well first of all you can write this as \[y=3 \ln(x-1) \cdot (x^2+1)^4 \]

myininaya (myininaya):

you need the product rule to continue

myininaya (myininaya):

\[y'=(3\ln(x-1))' \cdot (x^2+4)^4 + 3\ln(x-1) \cdot [(x^2+4)^4]'\]

myininaya (myininaya):

now continue with chain rule

OpenStudy (anonymous):

\[y=\ln \left[ \left( x-1 \right)^3\left( x^2+1 \right)^4 \right]=\ln \left( x-1 \right)^3+\ln \left( x^2+1 \right)^4\] \[=3\ln \left( x-1 \right)+4\ln \left( x^2+1 \right)\] \[\frac{ dy }{ dx }=\frac{ 3 }{ x-1 }+4*\frac{ 2x }{ x^2+1 }=?\]

myininaya (myininaya):

is that ln of that whole thing?

myininaya (myininaya):

ah maybe it is